OpenStudy (anonymous):

Can anyone answer this please? 3y to the 1/2 + y to the -1/2 Can anyone answer this please? 3y to the 1/2 + y to the -1/2 @Algebra

OpenStudy (asnaseer):

Assuming you mean:\[3y^{1/2}+y^{-1/2}=3y^{1/2}+\frac{1}{y^{1/2}}=\frac{3y+1}{y^{1/2}}\]